# Quantum Geometry: Revisiting electronic scales in quantum matter

### v1

@@ Abstract
Electronic properties of solids are often well understood via the low energy dispersion of Bloch bands, motivating single band approximations in many metals and semiconductors. However, a closer look reveals new length and time scales introduced by quantum dipole fluctuations due to interband mixing, which are reflected in the momentum space textures of the electronic wavefunctions. This structure is usually referred to as quantum geometry. The scales introduced by geometry not only qualitatively modify the linear and nonlinear responses of a material but can also have a vital role in determining the many-body ground state. This review explores how quantum geometry impacts properties of materials and outlines recent experimental advances that have begun to explore quantum geometric effects in various condensed matter platforms. We discuss the separation of scales that can allow us to estimate the significance of quantum geometry in various response functions.
I Introduction
The quantum many-body problem involves an exponentially large Hilbert space, making a full treatment impossible for the vast number of electrons typically found in a quantum material. Much of the analytic progress in the field has been achieved through low-energy theories that focus on a subset of degrees of freedom within a smaller, more manageable projected Hilbert space. While identifying this subspace is a challenge for many quantum materials, the first step is to determine the relevant energy, length, and time scales.
In a crystalline solid, the first relevant scale is the atomic lattice constant $a$, which allows for the definition of crystal momentum ${\bm{k}}$ that labels both band energies $\varepsilon({\bm{k}})$ and wavefunctions ${\psi({\bm{k}})}$. Much of the success in describing metals and semiconductors lies in the fact that electron transport can be described semiclassically, with the long wavelength behavior well captured by the band dispersion and its momentum derivatives such as the band velocity $v_{\bm{k}}\!\sim\!\partial_{\bm{k}}\varepsilon({\bm{k}})$ or band effective mass $1/m^{*}\!\sim\!\partial_{\bm{k}}^{2}\varepsilon({\bm{k}})$, combined with a phenomenological time scale which reflects how fast momentum relaxes in the presence of impurities or electron-electron interactions, $\tau$. Together with the Fermi velocity, it defines a mean free path of electrons $\ell_{\rm mr}=v_{F}\tau$, usually much larger than the lattice constant scale $a$.
In conventional condensed matter wisdom, the structure of materials at scales comparable with $a$ is usually considered unimportant to its collective behavior, justifying long-wavelength approximations such as effective field theories. Alternative lattice model approaches are simplified by taking a single site per unit cell, e.g., the Hubbard model, also reducing to a single band approximation. However, new insights from topology and quantum geometry say that, in fact, orbital mixing can lead to surprising qualitative changes in the physical behavior of materials. In this overview, we emphasize that orbital mixing across unit cells is the source of a new length scale that comes not from the band energies but by how electron wavefunctions change with momentum ${\partial_{\bm{k}}\psi({\bm{k}})}$. We refer to these quantities broadly as the quantum geometry of solids.
Quantum geometry in materials has a simple interpretation as the ground state dipole fluctuations [1]. This generalization to infinitely large systems follows from the well-understood case of dipole fluctuations in isolated atoms due to virtual level transitions [2]. Generally, these fluctuations can be used to characterize the size, shape, and angular momentum of atomic orbitals; however, in the lattice, qualitatively new orbitals can emerge from the quantum interference of distinct neighboring sites and are substantially richer than the isolated atomic problem. We argue that quantum geometry, with origin in virtual inter-band transitions can be used in solids to quantitatively describe the response of bound electrons in infinite lattices, even in metals where only part of the electrons are bound, while others are itinerant, plane-wave-like carriers. Strikingly, a quantum geometric description of bound electrons can be straightforwardly applied to systems without translation symmetry and beyond the single particle approximation. Therefore, it is a universal feature that systematically tracks the scale of dipole fluctuations.
As we overview here, dipole fluctuations have dramatic effects in linear and nonlinear responses of common materials, particularly those in low spatial dimensions such as graphene, transition metal dichalcogenides, or moiré heterostructures. Recent developments have also established that quantum geometry can play a fundamental role in the competition between correlated states, where the quantum geometry of the normal state can help stabilize exotic quantum orders such as fractionalized topological phases and superconductivity. In the following, we explain the concept of band geometry, focusing on dipolar fluctuations and their immediate consequences for the physical properties of quantum matter. We further discuss avenues to access the geometric tensor directly via different experimental probes and some of the intrinsic difficulties in their measurement.
@@ I Introduction
Figure 1: Dipole fluctuations lead to uncertainty in electron position. The typical spatial spread of localized electrons is characterized by a length scale $\ell_{g}$ determined by the quantum geometry of the ground state. a. The hydrogen atom, where the ground state wavefunction is an $s$ orbital with spread $\ell_{a}^{2}=3a_{B}^{2}/4$ where $a_{B}$ is the Bohr radius. b. Monolayer ${\rm MoTe}_{2}$, a semiconductor with $\sim 1\rm eV$ band gap. The localized electrons are shared between neighboring sites with a large spread spanning multiple unit cells due to nontrivial orbital interference.
II What is Quantum Geometry?
@@
Geometric phases lead to interference effects when states adiabatically traverse closed loops in parameter space. From the polarization of light [3], threading of a magnetic flux [4], or changes in atomic positions of molecules [5], geometric phases were understood to be general in quantum systems [6], arising from the Riemannian structure of the projected manifold of occupied quantum states. Provost and Vallee [7] introduced the quantum geometric tensor (QGT) to describe the adiabatic evolution in parameter space, showing that it is characterized not only by a phase rotation but also by a loss of the projected norm of the quantum states. Both effects are simultaneously characterized by the real and imaginary parts of the QGT. While Berry phase effects in electronic structure theory are now very well understood [8], the effects of the real part, known as the quantum metric, has remained mostly unexplored. Nonetheless, just as its imaginary counterpart, the real part is bound to play a significant role in quantum states of matter.
Fidelity susceptibility.
@@
The definition of the QGT is fairly universal and has only two requisites: a parameter space $\{\lambda_{\mu}\}$ and a projector $\hat{P}$, defining a subset of quantum states in the Hilbert space,
@@
where and $\partial_{\mu}\equiv\partial/\partial_{\lambda_{\mu}}$. We use the symbol without a hat $\mathcal{Q}_{\mu\nu}\!=\!\mathop{\mathrm{tr}}[\hat{\mathcal{Q}}_{\mu\nu}]$ to indicate the trace over the internal states of $\hat{P}$. Consider first the projector into a single state of an Hamiltonian $\mathcal{H}$, $\hat{P}=|\psi\rangle\langle\psi|$. When perturbations $\mathcal{H}+\sum_{\mu}\lambda_{\mu}\mathcal{H}_{\mu}^{\prime}$ are introduced, the parameters $\lambda_{\mu}$ define a parameter space and Eq. (1) its geometric tensor. In this case, the QGT captures the sensitivity of the state $\left|\psi\right>$ to perturbations, or fidelity susceptibility [9].
The fundamental reason behind quantum geometry is that the wavefunction overlap at adjacent points in parameter space deviates from unity, even when the wavefunctions are normalized at each point. Berry [6] famously demonstrated that wavefunctions acquire infinitesimal phases during this evolution, $\langle\psi(d\bm{\lambda})|\psi\rangle\approx 1-id\lambda_{\mu}\langle\psi|\partial_{\mu}\psi\rangle$, that accumulate to a gauge invariant geometric phase when the path is closed. The phase is captured by the Berry curvature $\Omega_{\mu\nu}=i\epsilon_{\mu\nu}\partial_{\mu}\langle\psi|\partial_{\nu}\psi\rangle$. A more subtle aspect is that the loss of norm is also gauge-invariant and measurable [10, 11]. Expanding the overlap to second order, we find
@@
where $g_{\mu\nu}$ can be seen as a metric that quantifies the distance between nearby quantum states. The QGT tensor proves invaluable in contexts where wavefunctions change abruptly, such as during quantum phase transitions [12], conical intersections of energy landscapes [5] or near singular band crossings in electric band structure [13], at which points the susceptibility Eq. (3) diverges [14, 15].
An elegant application of the geometric tensor for quantum materials was introduced by Walter Kohn [1], who studied a many-body ground state $\hat{P}_{0}=\left|\psi_{0}\right>\left<\psi_{0}\right|$ under twisted boundary conditions. The twist shifts the momentum in the spatial direction $\mu$ by the adiabatic parameter $\kappa_{\mu}$, $p_{\mu}\rightarrow p_{\mu}-\hbar\kappa_{\mu}$. Kohn argued that in the thermodynamic limit, the susceptibility Eq. (1) distinguishes the ground state of a metal from an insulator, diverging in metals while remaining finite for insulators [16]. This distinction based on the localization of wavefunctions bypasses the reference to a spectral gap, which can be ill-defined in disordered and many-body systems [1].
@@
The momentum shift introduced by Kohn can be interpreted as a susceptibility towards dipole transitions between states in $\hat{P}$ and states in the complementary projector $1-\hat{P}$. Since position and momentum are conjugate variables, $\partial_{\mu}\hat{P}\equiv i[\hat{r}_{\mu},\hat{P}]$, we may express Eq. (1) as
@@
where we have defined $\langle\cdot\rangle\!\equiv\!\mathop{\mathrm{tr}}[\hat{P}\cdot]$ as the expectation value at $T=0$. As it contains the dipole transitions between the ground state and excited states, an insulator can have nonzero quantum geometry as long as the ground state is not a position eigenstate. The diagonal elements of the real part of $\mathcal{Q}_{\mu\nu}$, i. e. the quantum metric $g_{\mu\nu}={\rm Re~{}}\mathcal{Q}_{\mu\nu}$ introduces a length scale $\ell_{g}=\sqrt{\mathop{\mathrm{tr}}g}$ associated with the size of dipole fluctuations in the system along the $\mu$ spatial direction. This fundamental insight was first discussed by Resta with the introduction of the “localization tensor” [17] and soon after by Souza, Wilkens and Martin [18].
@@
One does not need to look far to interpret the quantum metric as a measure of dipole fluctuations. A straightforward yet instructive example is an isolated Hydrogen atom subjected to an external electric field applied along $\hat{x}$. The perturbed Hamiltonian, $H^{\prime}=H-e\hat{x}E$, modifies the ground state wavefunction from $|\psi_{0}\rangle$ to $|\psi_{0}(E)\rangle$. The quantum metric is related to the fidelity via $g_{xx}=\partial|\langle\psi_{0}(E)|\psi_{0}\rangle|^{2}/\partial E^{2}$ as $E\rightarrow 0$. Using second-order perturbation theory, the metric becomes $g_{xx}=\sum_{m\neq 0}|\langle\psi_{0}|\hat{x}|\psi_{m}\rangle|^{2}=\langle(\hat{x}-\langle\hat{x}\rangle)^{2}\rangle_{0}$. This confirms that the metric measures the dipole fluctuations, which for the ground state of the hydrogen atom corresponds to the spatial spread of the $s$ atomic orbital, $\ell_{g}^{2}\!=\!(3/4)a_{B}^{2}$, with $a_{B}$ the Bohr radius. This example shows us that the size of the atomic orbital is a scale introduced by the geometric tensor in a single atom. Importantly, as long as there is time-reversal symmetry, the imaginary part of Eq. (3) vanishes exactly.
The advantage of Eq. (3) lies in its applicability to macroscopic systems where the behavior of electrons is qualitatively different from those in the Hydrogen atom. This is particularly important in bands with topological invariants where quantum geometry provides us with the language to capture the scale $\ell_{g}$ of dipole fluctuations of bound electrons, see Fig.1. The geometric length scale $\ell_{g}$ is usually comparable to the lattice constant $\ell_{g}\sim a$ and has dramatic signatures such as protected boundary modes [19] or quantization of Hall conductivity [20, 21]. For free electrons in a magnetic field, $\ell_{g}$ is nothing more than the magnetic length $\ell_{g}^{2}=\ell_{B}^{2}=\hbar/eB$, placing a similar formal ground to the quantum Hall effect in the continuum or in lattice models.
Here, it is our goal to emphasize that quantum geometry is indispensable in modeling materials with multiple bands - it presents a systematic way of including quantum corrections to the dipole operator beyond the conventional semi-classical approach. There has been increasing evidence that materials can exhibit drastically different properties solely due to differences in their quantum geometry [22]. This is particularly true when the band dispersion is significantly flattened by interference induce by, for example, a moiré potential. Therefore, the identification of observables capable of directly measuring or quantifying quantum geometry in real materials has emerged as an important challenge in the field [23].
III Quantum Geometry in Band Theory
@@
Figure 2: Kohn vs. single band metric. a. The Kohn metric uses the projected $\hat{P}^{0}$ and quantifies dipole transitions between occupied and unoccupied states. b. Quantum metric for a single isolated band with index $m$ using the band resolved projector $\hat{P}^{m}$ that includes all dipole transitions between band $m$ and all other bands $n\neq m$. The infinitesimal distance $|\langle u_{m,k}|u_{m,k+dk}\rangle|^{2}$ can be written as $1-g^{m}_{\mu\nu}dk_{\mu}dk_{\nu}$ where $g^{m}_{\mu\nu}=\sum_{n\neq m}\langle u_{m,k}|\hat{r}_{\mu}|u_{n,k}\rangle\langle u_{n,k}|\hat{r}_{\nu}|u_{m,k}\rangle$.
The localization tensor, defined in Eq. (3), quantifies the dipole matrix elements between the ground state and the excited states. Even in the case of non-interacting electrons, this formalism is often impractical because of the vast Hilbert space. However, from the simple case of an isolated Hydrogen atom, we can gain valuable intuition about the effects of dipole fluctuations and derive a geometric framework to apply in systems with multiple electrons. This approach comes with its challenges, as the definition of a dipole operator in an infinite periodic system has been settled only a few decades ago with the modern theory of polarization [24]. Central to this resolution were the cell-periodic parts of the Bloch wavefunction.
@@
Due to Bloch’s theorem, the electronic wavefunction in a periodic lattice potential can be decomposed into a plane-wave and a cell-periodic part, $|\psi_{n{\bm{k}}}\rangle=e^{i{\bm{k}}\cdot{\bm{R}}}|u_{n{\bm{k}}}\rangle$, where $n$ is the band index, ${\bm{k}}$ the crystal momentum and ${\bm{R}}$ a Bravais lattice vector. Being eigenstates of a Hermitian operator, the Bloch states are orthogonal. However, this orthogonality does extend to the cell-periodic function $|u_{n{\bm{k}}}\rangle$ which can have nonzero overlaps for different momenta. This introduces a notion of distance between two cell periodic states, giving rise to a Riemannian structure within the Brillouin Zone (BZ) by the adiabatic parameter ${\bm{k}}$ [25].
@@
To be more precise, we note that the quantum geometric tensor is the dipole correlation function that requires a projector. A natural choice is the projector to a Bloch state resolved in both band index and crystal momentum $\hat{P}_{n{\bm{k}}}$. However, it is important to notice that the dipole matrix element $\hat{r}_{\mu}^{mn}({\bm{k}},{\bm{k}}^{\prime})=\langle u_{m{\bm{k}}}|\hat{r}_{\mu}|u_{n{\bm{k}}^{\prime}}\rangle$ contains both intra- and interband contributions [26], (cf. Fig. 2),
@@
containing the matrix elements of the Berry connection $\left[\mathcal{A}_{\mu}({\bm{k}})\right]_{mn}=i\langle u_{m{\bm{k}}}|\partial_{\mu}u_{n{\bm{k}}}\rangle$, which are not by themselves gauge invariant. This does not pose an issue for QGT since the complementary projector in Eq. (3) removes the diagonal contributions of the position operator, making it gauge-independent. It can be compactly written in the band and momentum resolved basis as
$\mathcal{Q}_{\mu\nu}^{n}({\bm{k}})=\left\langle\partial_{\mu}u_{n{\bm{k}}}\right|1-\hat{P}_{n{\bm{k}}}\left|\partial_{\nu}u_{n{\bm{k}}}\right\rangle.$ (5)
@@
Berry curvature effects are well known in the semiclassical theory of transport in terms of the anomalous velocity [8]. Originally derived by Karplus and Luttinger through second-order perturbation theory [27], the anomalous velocity was later recognized as a consequence of wavefunction overlaps [8]. By explicitly deriving the equations of motion for a wave packet in the presence of an electric field, they demonstrated that the intrinsic anomalous velocity is precisely the Berry curvature [28, 29]. This paradigm shift has revolutionized our understanding of the Anomalous Hall effect, offering a fresh perspective distinct from conventional mechanisms based on scattering theory [30]. Furthermore, it has opened pathways to quantized Hall conductance by utilizing the quantization of the integral of Berry curvature over one band. This stems from the relation $\int_{\rm BZ}{\rm Im}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]=4\pi\mathcal{C}\epsilon_{\mu\nu}$ where $\int_{\rm BZ}\equiv\int d^{d}{\bm{k}}/(2\pi)^{d}$ and $\mathcal{C}$ is the Chern number that can only be an integer.

### v2

# Quantum Geometry and the Hidden Scales in Materials

@@ Abstract
Electronic properties of quantum materials solids are often well understood via the low energy dispersion of Bloch bands, motivating single band approximations in many metals and semiconductors. However, a closer look reveals length and time scales introduced by quantum dipole fluctuations due to interband mixing, which are reflected in the momentum space textures of the electronic wavefunctions. This structure is usually referred to as quantum geometry. These new scales not only qualitatively modify the linear and nonlinear responses of a material but can also have a vital role in determining the many-body ground state at low temperatures. In this Perspective, we explore how quantum geometry impacts properties of materials and outline recent experimental advances that have begun to explore quantum geometric effects in various condensed matter platforms. We discuss the separation of scales that can allow us to estimate the significance of quantum geometry in various response functions.
Introduction
Understanding the electronic properties of quantum materials relies on determining the relevant energy, length, and time scales to build a suitable model. In a crystalline solid, the fundamental length scale is the atomic lattice constant $a$, which allows for the definition of crystal momentum $k$ that labels both band energies $\varepsilon(k)$ and wavefunctions $\psi(k)$}. Much of the success in describing metals and semiconductors lies in the fact that electron transport can be described semiclassically. A phenomenological length scale, the electronic mean free path, captures the long wavelength behaviour. Because this mean-free path is typically much larger than the lattice constant, the structure of materials at the lattice scale $a$ is considered unimportant to the collective behavior at low-energies, justifying long-wavelength approximations and hydrodynamic descriptions. On the other end of the scale, attempting to correctly capture the effect of strong local interactions, microscopically motivated lattice models, like the Hubbard model, are often simplified by taking a single site per unit cell. In both of these regimes, a complicated quantum material problem is reduced to a single-band approximation, ignoring quantum interference effects across distinct orbits of the (often intricate) material unit cell.
However, orbital mixing can lead to surprising qualitative changes in the physical behavior of materials.The mixing within and across unit cells is the source of new length scales that come not from how band energies change, but by how electron wavefunctions change with crystal momentum $\partial_{k}\psi(k)$}. These changes are broadly referred to as the quantum geometry of solids. Apart from changes in the wavefunctions, a very convenient way to interpret quantum geometry in solids is as ground state dipole fluctuations [1], generalizing to infinitely large systems the well-understood case of dipole fluctuations in isolated atoms due to virtual level transitions [2]. Generally, these fluctuations can be used to characterize the size, shape, and angular momentum of atomic orbitals. Strikingly, on the lattice, qualitatively new orbitals can emerge from the quantum interference of distinct neighboring sites and are substantially richer than the isolated atomic problem. Dipole fluctuations have dramatic effects in linear and nonlinear responses of common materials, particularly those in low spatial dimensions such as graphene, transition metal dichalcogenides, or moiré heterostructures.
Over the last few years, with the rapid advances in the field of two-dimensional heterostructures, there has been mounting evidence that quantum materials can exhibit distinct transport properties, such as superfluid or exciton stiffness, solely due to differences in their quantum geometry [3, 4]. This effect is widespread and can be appreciated by a simple example: rock salt and diamond reflect light very differently. Their energy gaps are comparable, but their electronic wavefunctions are remarkably different. The effects of wavefunction geometry on the response of materials can be particularly striking in flat-band systems, where they can qualitatively change the physical properties. Such flat bands can now be achieved by moiré engineering in a variety of materials [5, 6, 7]. Therefore, identifying observables capable of directly measuring or quantifying quantum geometry in materials has become an important goal in the field. Quantum geometry can also play a fundamental role in the competition between correlated states, where the quantum geometry of the normal state can help stabilize exotic quantum orders such as fractionalized topological phases and superconductivity.
In this Perspective, it is our goal to emphasize that quantum geometry is an essential feature of any model with multiple bands. We aim to highlight that quantum geometry can be used in solids to quantitatively and systematically describe the response of bound electrons in infinite lattices. As we argue here, this framework is valid even for metals where only part of the electric charge is bound, coexisting with itinerant plane-wave electron states. Furthermore, the quantum geometric description of bound electrons is straightforwardly applied to systems without translation symmetry and beyond the single particle approximation. Therefore, it is a universal feature that systematically tracks the length and time scales that capture dipole dynamics in solids. In the following, we explain the concept of band geometry, focusing on its real space interpretation, and its immediate consequences for the physical properties of quantum matter. We further discuss possible experimental probes to measure the geometric tensor and highlight some of the intrinsic difficulties in their measurement.
@@ Introduction
Figure 1: Dipole fluctuations lead to uncertainty in electron position. The corresponding spatial spread is characterized by the length scale $\ell_{g}$ determined by the quantum geometry of the ground state. a. The hydrogen atom, where the ground state wavefunction is an $s$ orbital with spread $\ell_{a}^{2}=3a_{B}^{2}/4$ where $a_{B}$ is the Bohr radius. b. Monolayer ${\rm MoTe}_{2}$, a semiconductor with $\approx 1\rm eV$ band gap and a topological obtruction. The localized electrons are shared between neighboring sites with a large spread spanning multiple unit cells due to nontrivial orbital interference. The geometric scale $\ell_{g}$ is comparable to the lattice constant.
What is Quantum Geometry?
@@
When a quantum state adiabatically moves through a parameter space to make a closed loop, interference effects can cause the state to pick up a phase difference, called the geometric phase. This can take many forms, from the polarization of light [8], threading of a magnetic flux [9], or changes in atomic positions of molecules [10]. Geometric phases are understood to be general in quantum systems [11], arising from the Riemannian structure of the projected manifold of occupied quantum states. The quantum geometric tensor (QGT)[12] can be used to describe the adiabatic evolution in parameter space, showing that it is characterized not only by a phase rotation but also by a loss of the projected norm of the quantum states. Both effects are simultaneously characterized by the real and imaginary parts of the QGT. The imaginary part of the QGT, which describes the Berry curvature, or Berry phase, is well understood in electronic structure theory [13]. However, the effects of the real part, the quantum metric, has remained mostly unexplored. Nonetheless, just as its imaginary counterpart, the real part is bound to play a significant role in quantum states of matter.
Fidelity susceptibility
@@
The definition of the QGT is universal and has only two requisites: a parameter space $\{\lambda_{\mu}\}$ and a projector $\hat{P}$, defining a subset of quantum states in the Hilbert space,
@@
where $\partial_{\mu}\equiv\partial/\partial_{\lambda_{\mu}}$. We use the symbol without a hat $\mathcal{Q}_{\mu\nu}\!=\!\mathop{\mathrm{tr}}[\hat{\mathcal{Q}}_{\mu\nu}]$ to indicate the trace over the internal states of $\hat{P}$. Consider the projector into a single state of an Hamiltonian $\mathcal{H}$, $\hat{P}=|\psi\rangle\langle\psi|$. When perturbations $\mathcal{H}+\sum_{\mu}\lambda_{\mu}\mathcal{H}_{\mu}^{\prime}$ are introduced, the parameters $\lambda_{\mu}$ define a parameter space and Eq. (1) its geometric tensor. In this case, the QGT captures the sensitivity of the state $\left|\psi\right$> to perturbations, or fidelity susceptibility [14]. The fundamental reason behind quantum geometry is that the wavefunction overlap at adjacent points in parameter space deviates from unity, even when the wavefunctions are normalized at each point. Michael Berry [11] demonstrated that wavefunctions acquire infinitesimal phases during this evolution, $\langle\psi(d\bm{\lambda})|\psi\rangle\approx 1-id\lambda_{\mu}\langle\psi|\partial_{\mu}\psi\rangle$, that accumulate to a gauge-invariant geometric phase when the path is closed. The phase is captured by the Berry curvature $\Omega_{\mu\nu}=i\epsilon_{\mu\nu}\partial_{\mu}\langle\psi|\partial_{\nu}\psi\rangle$. Concurrently, the reduction in magnitude of the wavefunction overlap is also gauge-invariant [15, 16]. Expanding the wavefunction overlap to second order, we find
@@
where $g_{\mu\nu}$ is the quantum metric that quantifies the distance between nearby quantum states. The QGT tensor signals where wavefunctions change abruptly, such as during quantum phase transitions [17], conical intersections of energy landscapes [10] or near singular band crossings in electronic band structure [18], at which points the susceptibility (Eq. (1)) diverges [19, 20].
An elegant application of the geometric tensor for quantum materials was introduced by Walter Kohn in the 1960s [21], who studied a many-body ground state $\hat{P}^{0}=\left|\psi_{0}\right>\left<\psi_{0}\right|$ under twisted boundary conditions. The twist shifts the momentum in the spatial direction $\mu$ by the adiabatic parameter $\kappa_{\mu}$, $p_{\mu}\rightarrow p_{\mu}-\hbar\kappa_{\mu}$. Kohn argued that in the thermodynamic limit, the susceptibility, as given by the QGT (Eq. (1)) distinguishes the ground state of a metal from an insulator, diverging in metals while remaining finite for insulators [22]. This distinction based on the localization of wavefunctions bypasses the reference to a spectral gap, which can be ill-defined in disordered and many-body systems.
@@
The momentum shift under twisted boundary conditions can be interpreted as a susceptibility towards dipole transitions between states in $\hat{P}$ and states in the complementary projector $1-\hat{P}$. Since position and momentum are conjugate variables, $\partial_{\mu}\hat{P}\equiv i[\hat{r}_{\mu},\hat{P}$], we may express Eq. (1) as a localization tensor[23, 24]:
@@
where we have defined $\langle\cdot\rangle\!\equiv\!\mathop{\mathrm{tr}}[\hat{P}\cdot$] as the expectation value at $T=0$. As the localization tensor contains the dipole transitions between the ground state and excited states, an insulator will have nonzero quantum geometry as long as the ground state is fully localized, and not a position eigenstate. The quantum metric, $g_{\mu\nu}={\rm Re~}\mathcal{Q}_{\mu\nu}$, introduces a length scale $\ell_{g}=\sqrt{\mathop{\mathrm{tr}}g}$ associated with the size of dipole fluctuations.
@@
A straightforward yet instructive example to view the quantum metric as a measure of dipole fluctuations is to consider an isolated hydrogen atom subjected to an external electric field. The perturbed Hamiltonian, $H^{\prime}=H-e\hat{x}E$ (for a field along $\hat{x}$) modifies the ground state wavefunction from $|\psi_{0}\rangle$ to $|\psi_{0}(E)\rangle$. The quantum metric is related to the fidelity via $g_{xx}=\partial|\langle\psi_{0}(E)|\psi_{0}\rangle|^{2}/\partial E^{2}$ as $E\rightarrow 0$. Using second-order perturbation theory, the metric becomes $g_{xx}=\sum_{m\neq 0}|\langle\psi_{0}|\hat{x}|\psi_{m}\rangle|^{2}=\langle(\hat{x}-\langle\hat{x}\rangle)^{2}\rangle_{0}$. Therefore the metric measures the dipole fluctuations, which for the ground state of the hydrogen atom corresponds to the spatial spread of the $s$ atomic orbital, $\ell_{g}^{2}\!=\!(3/4)a_{B}^{2}$, with $a_{B}$ the Bohr radius. This example shows us that in a single atom, the size of the atomic orbital is a scale introduced by the geometric tensor.
The localization tensor (Eq. (3)) is useful in that it can be applied to macroscopic systems where the behavior of electrons is qualitatively different from those in the hydrogen atom. Quantum geometry provides us with a language to capture the scale $\ell_{g}$ of dipole fluctuations of bound electrons in materials in infinite lattices, whose wavefunctions in the clean limit are fully delocalized Bloch states (Fig.1). The geometric length scale $\ell_{g}$ is typically on the order of the lattice constant ($\ell_{g}\approx a$). However, there are important exceptions where $\ell_{g}\gg a$. In a massive Dirac system, for instance, $\ell_{g}$ diverges as the energy gap closes and can remain much larger than the lattice constant for a small mass. This behaviour underscores that $\ell_{g}$ arises from wavefunction interference across many sites and is not solely determined by local energetics.
In topological systems, such as those exhibiting quantized Hall conductivity [25, 26] or symmetry-protected boundary modes [27], topological invariants place a generic lower bound on the magnitude of dipole fluctuations. This bound becomes saturated in idealized scenarios, such as free electrons in a magnetic field where the geometric scale coincides with the magnetic length $\ell_{g}^{2}=\ell_{B}^{2}=\hbar/eB$ [28]. In trivial systems with weakly hybridized orbitals, the geometric scale captures the sizes of the orbitals themselves $\ell_{g}\approx a_{B}$, and can often be ignored, as is the case in single band approximations. Crucially, the scale $\ell_{g}$ is universally a well-defined indicator of dipole fluctuations.
@@
Figure 2: Choice of projected manifold in defining quantum metric. a. The Kohn metric is defined using the ground-state projector $\hat{P}^{0}$ and captures dipole transitions between occupied and unoccupied states. b. The quantum metric for a single isolated band $m$ is defined using the band-resolved projector $\hat{P}^{m}$, incorporating dipole transitions between band $m$ and all other bands $n\neq m$. The infinitesimal overlap between Bloch states is given by $|\langle u_{m,\bm{k}}|u_{m,\bm{k}+d\bm{k}}\rangle|^{2}=1-g^{m}_{\mu\nu}\,dk_{\mu}\,dk_{\nu}$, where $g^{m}_{\mu\nu}=\sum_{n\neq m}\langle u_{m,\bm{k}}|\hat{r}_{\mu}|u_{n,\bm{k}}\rangle\langle u_{n,\bm{k}}|\hat{r}_{\nu}|u_{m,\bm{k}}\rangle$. The Kohn metric appears in optical sum rules, while the single-band metric governs the superfluid stiffness in flat-band systems. The two are formally distinct, in that they differ in their choice of the projected manifold, and are thus not interchangeable.
Quantum Geometry in Band Theory
@@
The localization tensor (Eq. (3)), quantifies the dipole matrix elements between the ground state and the excited states. Even in the case of non-interacting electrons, this formalism is often impractical because of the vast Hilbert space. However, from the simple case of an isolated hydrogen atom, we can gain valuable intuition about the effects of dipole fluctuations and derive a geometric framework to apply in systems with multiple electrons. This approach requires a careful definition of the dipole operator in an infinite periodic system, which is provided by the modern theory of polarization [29]. Due to Bloch’s theorem, the electronic wavefunction in a periodic lattice potential can be decomposed into a plane-wave and a cell-periodic part, $|\psi_{n{\bm{k}}}\rangle=e^{i{\bm{k}}\cdot\hat{{\bm{r}}}}|u_{n{\bm{k}}}\rangle$, where $n$ is the band index, $\bm{k}$} the crystal momentum and $\hat{\bm{r}}$ the position operator. Being eigenstates of a Hermitian operator, the Bloch states are orthogonal. However, this orthogonality does extend to the cell-periodic function $|u_{n{\bm{k}}}\rangle$ which can have nonzero overlaps for different momenta. This introduces a notion of distance between two cell periodic states, giving rise to a Riemannian structure within the Brillouin Zone (BZ) by the adiabatic parameter $\bm{k}$} [30].
@@
To be more precise, we note that the quantum geometric tensor is the dipole correlation function that requires a projector. A natural choice is the projector to a Bloch state resolved in both band index and crystal momentum $\hat{P}_{{\bm{k}}}^{n}$. However, it is important to notice that the dipole matrix element $\hat{r}_{\mu}^{mn}({\bm{k}},{\bm{k}}^{\prime})=\langle u_{m{\bm{k}}}|\hat{r}_{\mu}|u_{n{\bm{k}}^{\prime}}\rangle$ contains both intra- and interband contributions [31] (Fig. 2),
@@
containing the matrix elements of the Berry connection $\left[\mathcal{A}_{\mu}({\bm{k}})\right]_{mn}=i\langle u_{m{\bm{k}}}|\partial_{\mu}u_{n{\bm{k}}}\rangle$, which are not by themselves gauge invariant. A crucial aspect of the quantum geometric tensor is that the complementary projector in Eq. (3) removes the diagonal contributions of the position operator, and therefore, it is fully gauge independent and measurable. For Bloch electrons, the QGT can be compactly written in the band and momentum basis as
$\mathcal{Q}_{\mu\nu}^{n}({\bm{k}})=\left\langle\partial_{\mu}u_{n{\bm{k}}}\right|1-\hat{P}_{{\bm{k}}}^{n}\left|\partial_{\nu}u_{n{\bm{k}}}\right\rangle.$ (5)
@@
Berry curvature effects are well known in the semiclassical theory of transport in terms of the anomalous velocity [13]. Originally derived through second-order perturbation theory [32], the anomalous velocity was later recognized as a consequence of wavefunction overlaps. By explicitly deriving the equations of motion for a wave packet in the presence of an electric field, it has been demonstrated that the intrinsic anomalous velocity is precisely the Berry curvature [33, 34]. Understanding the connection between the anomalous velocity and wavefunction overlap has revolutionized our understanding of the Anomalous Hall effect, offering a fresh perspective distinct from conventional mechanisms based on scattering theory [35]. Furthermore, it has opened pathways to quantized Hall conductance by utilizing the quantization of the integral of Berry curvature over one band. This stems from the relation $\int_{\rm BZ}{\rm Im}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]=4\pi\mathcal{C}\epsilon_{\mu\nu}$ where $\int_{\rm BZ}\equiv\int d^{d}{\bm{k}}/(2\pi)^{d}$ and $\mathcal{C}$ is the Chern number that can only be an integer.